Shift report 20030115 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Jan 15 2003 - 12:06:49 EST

  • Next message: Apache: "Shift report 20030116 16:00-24:00"
    ------------------------------------------------------------
    Date: Wed Jan 15 12:06:49 2003, Shift: 08:00-16:00
    Supervisor: ykl, with: 
    Log book#13 pages 64 - 64
    
    Shift Summary:
    ------------------------------------------------------------
    1. RHIC was in beam study mode with widely fluctuation
    background (H1 and H2 going over 1 megaherz occasionally).
    
    2. No useable beam for data. TPC's, DVM and DC's were turned
    off (not just parked) during most of the shift since the
    background was too high.
    
    DB replaced discriminators in the trigger system (TD1, INL2,
    TOFW).
    
    3. In H2, H2101 and H2103 showed zero measured HV although
    demand voltage was 1600V for both. 
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Wed Jan 15 2003 - 12:08:20 EST